OpenStudy (mendicant_bias):

Now dealing with Dirac-Delta functions in ODE and their laplaces/inverses. Posted below momentarily.

OpenStudy (mendicant_bias):

\[y'-3y=\delta(t-2), \ \ \ y(0)=0\]

OpenStudy (mendicant_bias):

Taking the Laplace of both sides, \[sY(s)-y(0)-3Y(s)=e^{-2s}\]

OpenStudy (sidsiddhartha):

u got it

OpenStudy (mendicant_bias):

\[Y(s)=\frac{e^{-2s}}{s-3}\]

OpenStudy (mendicant_bias):

\[f(t)=\mathcal{L^{-1}}\left\{\vphantom{}\frac{e^{-2s}}{s-3}\right\}= \ ?\] (Haven't dealt with inverse laplaces with e^(something)'s in their argument for a while, lookuing up how to do so)

OpenStudy (mendicant_bias):

It's some sort of shifting, right? Shifting on the t-axis?

OpenStudy (sidsiddhartha):

yes same thing

OpenStudy (mendicant_bias):

Alright,\[e^{3t}\mathcal{U}(t-2)\]I think

OpenStudy (mendicant_bias):

Whoops, 3t-2

OpenStudy (sidsiddhartha):

\[L\frac{ 1 }{ s-3 }=e^{3t}\]

OpenStudy (mendicant_bias):

3(t-2)

OpenStudy (mendicant_bias):

\[e^{3(t-2)}\mathcal{U}(t-2)\]

OpenStudy (sidsiddhartha):

now just make it lag by -2

OpenStudy (mendicant_bias):

yeah, think I got it (not sure if my posts are lagging)

OpenStudy (sidsiddhartha):

yeah that looks fine

OpenStudy (mendicant_bias):

Alright! Continual progress!!!!!

OpenStudy (sidsiddhartha):

yes same here lagging

ganeshie8 (ganeshie8):

ur replies are dancing here

OpenStudy (mendicant_bias):

Gonna keep going, don't let me keep you @sidsiddhartha , but thank you, again. Heh, I saw mine doing the same earlier, yeah

OpenStudy (sidsiddhartha):

lol

OpenStudy (sidsiddhartha):

i can stay a little longer :)